HUBBLE SPACE TELESCOPE DAILY REPORT #2532 PERIOD COVERED: 0000Z (UTC) 01/04/00 - 0000Z (UTC) 01/05/00 Daily Status Report as of 005/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8530 (SMOV3a STIS BEA Science) The Space Telescope Imaging Spectrograph (CCD) was used to obtain SMOV3a Bright Earth Avoidance time period for science exposures with the STIS CCD. Various kinds of targets are observed. The proposal completed nominally. 1.2 Completed Two Sets of WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 1.3 Completed S/C 8518 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} is used to characterize the HST dynamic response due to environmental disturbances during extended gyro hold intervals. Disturbances are caused by on-board mechanism motion and transient effects resulting from thermal strain. These disturbances excite the natural modes of HST structural appendages and produce vehicle dynamic response as measured by the rate gyros. Disturbance signatures are compared with signatures taken after previous servicing missions to identify changes in mechanisms or structures since the last VDT. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. At least 5 orbits at +V3 sunpoint {with 90-degree solar array angle} and 3 orbits at -V1 sunpoint {with 0-degree solar array angle} are required for this test. Using two different solar array orientations allows additional identification of disturbances induced by the solar arrays. Low bandwidth attitude control gains are used to configure the HST for "nearly open-loop" dynamic response during the test. This configuration allows observation of HST disturbance response that is nearly independent of the affects of the pointing control law. No science instruments are used during this test. The test completed with no reported problems. 1.4 Completed WF/PC-2 8497 (Point Spread Function Verification) The WF/PC-2 was used to verify the WFPC2 PSF after SM3A. There were no reported problems. 1.5 Completed STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.6 Completed Two Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. The observations completed with no reported problems. 1.7 Completed Two Sets of STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.8 Completed STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 15 Successful: 15 Scheduled Re-acquisitions: 1 Successful: 1 2.2 FHST Updates: Scheduled: 31 Vehicle Disturbance Test Successful: 31 2.3 Operations Notes: Servicing Mission Orbital Verification (SMOV) activities are being performed. The Vehicle Disturbance Test (VDT) installation procedure was performed at 004/1520Z per an operations request. The test ran nominally as vehicle rates, attitude response, and solar array disturbances were monitored. The next iteration of the test will occur early on day 007. ROP SR-8 was utilized at 004/1521Zto configure the AER start and stop commands. A NICMOS memory dump was commanded at 004/164556Z. For reasons that are not currently understood, the CCS system did not receive the dump, although all required directives and procedures were followed. HSTAR 7461 was written. Later, at 004/210033Z, the NICMOS dump was successfully performed. The SSA transmitter was turned on and off at 004/1644Z as directed by ROP IC-2. There was a STIS EMC retry at 005/015116Z. HSTAR 7462 was written. ROP NS-12 was used to clear the STIS flight software error counter at 005/0219Z. A TTR was written when the GCMR capability was lost at 005/0801Z during a TDE event. As a result, there was a 2-minute, 10-second non-recoverable loss of engineering data. ESTR-2 was reconditioned at 005/0840Z, using ROP RD-7. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW